# A path-dependent PDE solver based on signature kernels

### v3

@@ 2.3.3. Options under rough Bergomi model
(2.9) $\left\{\begin{array}[]{rl}&\partial_{t}u+\frac{1}{2}\psi_{t}(\gamma_{t})\Big{(}\partial_{x}^{2}u-\partial_{x}u\Big{)}+\frac{1}{2}\widetilde{\partial}_{\gamma}^{K^{t}K^{t}}u+\rho\sqrt{\psi_{t}(\gamma_{t})}\widetilde{\partial}_{\gamma}^{K^{t}}(\partial_{x}u)=0,\\ &u(T,x,\gamma)=\phi(x).\end{array}\right.$
@@ 2.3.4. An important remark on the singularity of the kernel
We emphasise that the PPDEs presented in these examples fit in (2.3) albeit with a non-continuous direction $K^{t}$. The pathwise derivative is thus defined as the limit of the regularised derivative with the approximation $K^{t,\delta}$ as in (2.2). One can show thanks to the probabilistic representations that the solution to the regularised PPDE converges to the solution of the singular one as $\delta$ goes to zero. Our numerical scheme approximates the PPDE (2.3) with a regularised kernel, for a fixed $\delta>0$.
@@ 2.4. Functions on bounded variation paths
Making the framework of subsection 2.2 amenable to the application of signature kernels requires some carefulness. Firstly, the latter are defined on bounded variation paths, a subset of the continuous paths considered in the previous subsection. In this subsection, functions are thus defined over $\Omega$ instead of $\widetilde{\Omega}$.
Further, we make the observation that all three examples presented above are functionals $(t,x,\gamma)\mapsto u(t,x,\gamma)$ that, for all $t\in\mathbb{T}$, only act on $(\gamma_{s})_{s\in[t,T]}$, the path after time $t$. This crucial property is the reason why we want to consider paths that are constant on $[0,t]$, so that no unnecessary information is included. We denote by $\mathcal{C}(\Omega)$ the set of continuous functions $u:\Omega\to\mathbb{R}$ with respect to $\bm{d}$; for those there exists a family $(u_{t,x})_{t\in\mathbb{T},x\in\mathbb{R}^{d}}$ of continuous maps from $\mathrm{BV}_{t}$ to $\mathbb{R}$ such that $u(t,x,\gamma)=u_{t,x}(\gamma^{0}\lvert_{[t,T]},\check{\gamma}\lvert_{[t,T]})$. Note that this distance looks into the path on $[0,t]$, indeed without this information one would lose the positivity for two paths distinct on $[0,t]$, while considering only continuity of $u_{t,x}$ leads to issues in defining compact sets of $\Omega$. To be consistent with the derivative of the signature kernel introduced in [47], we choose a weaker form of pathwise derivative than the Fréchet one: the Gateaux derivative (it will appear later that signatures are continuous with respect to the $\mathcal{C}_{0}$ norm and therefore both derivatives agree). This is where the restriction to $[t,T]$ becomes useful and allows us to avoid jumps. Recall that the Gateaux derivative perturbs the path on the whole interval where it is defined
@@
We instead define the pathwise derivative of $u\in\mathcal{C}(\Omega)$ in the direction $\eta\in\mathrm{BV}_{0}$ such that only the path $\check{\gamma}$ is perturbed and solely on $[0,t]$:
(2.11) ^$\widehat{\partial}^{\eta}_{\gamma}u(t,x,\gamma):=\partial^{({\bm{0}},\check{\eta}\lvert_{[t,T]})}_{\gamma}u_{t,x}(\gamma^{0}\lvert_{[t,T]},\check{\gamma}\lvert_{[t,T]})=\lim_{\varepsilon\to 0}\frac{u_{t,x}(\gamma^{0}\lvert_{[t,T]},(\check{\gamma}+\varepsilon\check{\eta})\lvert_{[t,T]})-u_{t,x}(\gamma^{0}\lvert_{[t,T]},\check{\gamma}\lvert_{[t,T]})}{\varepsilon}.$
Note that this amounts to three different notions of pathwise derivatives $\widetilde{\partial},\widehat{\partial},\partial$. Similarly to the former setup in subsection 2.2, for $n,n_{1},n_{2},n_{3}\in\mathbb{N}_{0}$, we can define the higher derivatives $\partial_{t}^{n_{1}},\,\partial_{x}^{n_{2}}$ and ^$\widehat{\partial}_{\gamma}^{\bm{\eta}}$, for $\bm{\eta}=(\eta_{1},\cdots,\eta_{n_{3}})\in(\mathrm{BV}_{0})^{n_{3}}$. We denote $\mathcal{C}^{\mathcal{N}}(\Omega)$ the space of functions admitting continuous such pathwise derivatives of order $\mathcal{N}\subset\mathbb{N}^{3}_{0}$. We can now define PPDEs on $\Omega$ similar to (2.3)
(2.12) $\left\{\begin{array}[]{rl}&\mathcal{L}_{t}u(t,x,\gamma)=g(t,x,\gamma),\\ &u(T,x,\gamma)=\phi(T,x,\gamma),\end{array}\right.$
where $f,\phi:\Omega\to\mathbb{R}$ are measurable functions and the linear operator $\mathcal{L}_{t}$ is a combination of derivatives of all arguments, analogue of ~$\widetilde{\mathcal{L}}_{t}$ for functions independent of $\gamma_{[0,t]}$, albeit with a pathwise derivative of Gateaux type instead of Fréchet. More precisely, there exist $\mathcal{N}\subset\mathbb{N}^{3}_{0}$, $a_{n}\in\mathbb{R}$ for all $n\in\mathcal{N}$ and $\bm{\eta}\in(\mathrm{BV}_{0})^{n_{3}}$ such that
@@ 2.5. A kernel method for PPDEs
We saw that we cannot solve a PPDE of the type ~$\widetilde{\mathcal{L}}_{t}u=f$ on $\widetilde{\Omega}$ directly with signature kernel methods, hence why we defined an analogous weaker PPDE $\mathcal{L}_{t}u=f$ on $\Omega\subset\widetilde{\Omega}$ which solution can be suitably represented with signature kernels. Functions admitting Fréchet derivatives also admit Gateaux derivatives and they are equal, therefore solutions to the former PPDE are also solutions to the latter. This will allow us in Theorem 4.4 to bridge the gap between the equation we solve and the equation we target. Now that we have this picture in mind, we can present the gist of our method, inspired by [12] and extended to the path-dependent case.
We consider a kernel $\kappa$, indexed on a compact subset $\mathcal{X}$ of $\Omega$, as product of an RBF kernel and a signature kernel indexed on $\mathbb{R}^{d+1}$ and $\mathrm{BV}_{0}$ respectively. We call $\mathcal{H}$ its associated RKHS. We pick $m+n$ collocation points $\omega^{1},...,\omega^{m+n}\in\mathcal{X}$ of the form $\omega^{i}=(t^{i},x^{i},\gamma^{i})$ and order them in such a way that $0\leq t^{i}<T$ for $i=1,...,m$ and $t^{i}=T$ for $i=m+1,...,m+n$. The former belong to the interior of the domain while the latter correspond to the boundary (terminal) condition. The main idea consists in approximating the solution of the PPDE (2.12) with a minimiser of the following optimal recovery problem
@@
In other words, we search for the function with minimal RKHS norm that satisfies the PDE constraints at all the collocation points. This problem formulation raises a number of questions, of both theoretical and practical importance:
@@
    Does the domain of $\mathcal{L}_{t}$ contain $\mathcal{H}$
@@
    Does the PPDE (2.3) have a unique classical solution $u^{\star}$
@@
    Does $u_{m,n}$ converge to $u^{\star}$ as $m,n\to\infty$
After designing our kernel and RKHS in Section 3, we answer positively to all these questions under appropriate conditions. In order to successively execute limit arguments, we exploit that $\mathcal{X}$ is a compact set of $\Omega$, which we define explicitly under the topology induced by $\bm{d}$. Exploiting the robust signatures introduced in [14] may allow to lift this assumption; we will investigate this idea in the future. The interested reader may find a summary of the answers as follows.
@@
    The optimal recovery problem reduces to a finite dimensional quadratic optimisation problem with linear constraints. The latter can be solved via gradient descent, aloof from sophisticated learning algorithms. See Section 4.1.
@@
    Theorem 4.4 ensures convergence towards the true solution under two different sets of assumptions, both requiring that $u^{\star}$ is the unique solution of the PPDE (2.3) and that the collocation points form a countable dense set of $\mathcal{H}$.
@@
Finally, let us summarise the several levels of approximation. We numerically solve the optimisation problem (2.14), which is a discretisation of the PPDE (2.12) over $\Omega$. Under the right set of assumptions, the sequence of solutions to the optimisation problem converges the solution to the PPDE (2.3) over $\widetilde{\Omega}$. Finally, the latter is an approximation of the PPDE with a singular direction $\bm{\eta}$. While our setup is designed to match PPDEs of second generation (arising from Volterra processes), it could be adapted to the first generation of Dupire (corresponding to path-dependent payoffs).

### v4

@@ 2.3.3. Options under rough Bergomi model
(2.12) $\left\{\hskip-11.38092pt\begin{array}[]{rl}&\partial_{t}u+\frac{1}{2}\psi_{t}(\gamma_{t})\Big(\partial_{x}^{2}u-\partial_{x}u\Big)+\frac{1}{2}\widetilde{\partial}_{\gamma}^{K^{t}K^{t}}u+\rho\sqrt{\psi_{t}(\gamma_{t})}\widetilde{\partial}_{\gamma}^{K^{t}}(\partial_{x}u)=0,\\ &u(T,x,\gamma)=\phi(x).\end{array}\right$.
@@ 2.3.4. An important remark on the singularity of the kernel
We emphasise that the PPDEs presented in these examples fit in (2.6) albeit with a non-continuous direction $K^{t}$. Since the pathwise derivative in this direction is defined as the limit of the regularised derivative with direction $K^{t,\delta}$ as in (2.5), one can show thanks to the probabilistic representations that the solution to the regularised PPDE converges to the solution of the singular one as $\delta$ goes to zero. Our numerical scheme approximates the PPDE (2.6) with a regularised kernel, for a fixed $\delta>0$.
@@ 2.4. Functions on bounded variation paths
Making the framework of subsection 2.2 amenable to the application of signature kernels requires some care. Firstly, the latter are defined on bounded variation paths, a subset of the continuous paths considered in the previous subsection. In this subsection, functions are thus defined over $\Omega$ (2.2) instead of $\widetilde{\Omega}$ (2.1).
Further, we make the observation that all three examples presented above are functionals $(t,x,\gamma)\mapsto u(t,x,\gamma)$ that, for all $t\in\mathbb{T}$, only act on $(\gamma_{s})_{s\in[t,T]}$, the path after time $t$. This crucial property is the reason why we want to consider paths that are constant on $[0,t]$, so that no unnecessary information is included. We denote by $\mathcal{C}(\Omega)$ the set of continuous functions $u:\Omega\to\mathbb{R}$ with respect to $\bm{d}$ (2.3); for those there exists a family $(u_{t,x})_{t\in\mathbb{T},x\in\mathbb{R}^{d}}$ of continuous maps from $\mathrm{BV}_{t}=\mathcal{C}^{1-var}([t,T],\mathbb{R}^{e})$ to $\mathbb{R}$ such that $u(t,x,\gamma)=u_{t,x}(\gamma^{0}\lvert_{[t,T]},\check{\gamma}\lvert_{[t,T]})$ (recall that for each time-augmented path $\gamma$ we denote by $\gamma^{0}$ the strictly monotone coordinate and by $\check{\gamma}$ the $\mathbb{R}^{e-1}$-valued path without the strictly monotone coordinate). Note that the distance $\bm{d}$ looks into the path on $[0,t]$, indeed without this information one would lose the positivity for two paths distinct on $[0,t]$, while considering only continuity of $u_{t,x}$ leads to issues in defining compact sets of $\Omega$. To be consistent with the derivative of the signature kernel introduced in [49], we choose a weaker form of pathwise derivative than the Fréchet one: the Gateaux derivative (it will appear later that signatures are continuous with respect to the supremum norm and therefore both derivatives agree). This is where the restriction to $[t,T]$ becomes useful and allows us to avoid jumps. Recall that the Gateaux derivative perturbs the path on the whole interval where it is defined
@@
We instead define the pathwise derivative of $u\in\mathcal{C}(\Omega)$ in the direction $\eta\in\mathrm{BV}_{0}$ such that only the path $\check{\gamma}$ is perturbed and solely on $[t,T]$:
(2.14) $\widehat{\partial}^{\eta}_{\gamma}u(t,x,\gamma):=\partial^{({\bm{0}},\check{\eta}\lvert_{[t,T]})}_{\gamma}u_{t,x}(\gamma^{0}\lvert_{[t,T]},\check{\gamma}\lvert_{[t,T]})=\lim_{\varepsilon\to 0}\frac{u_{t,x}(\gamma^{0}\lvert_{[t,T]},(\check{\gamma}+\varepsilon\check{\eta})\lvert_{[t,T]})-u_{t,x}(\gamma^{0}\lvert_{[t,T]},\check{\gamma}\lvert_{[t,T]})}{\varepsilon}.$
In contrast to the Fréchet derivative defined in (2.4) that perturbs the path over the interval $[t,T]$, this definition deals only with continuous (even bounded variation) paths. Similarly to the former setup in subsection 2.2, for $n,n_{1},n_{2},n_{3}\in\mathbb{N}_{0}$, we can define the higher derivatives $\partial_{t}^{n_{1}},\,\partial_{x}^{n_{2}}$ and $\widehat{\partial}_{\gamma}^{\bm{\eta}}$, for $\bm{\eta}=(\eta_{1},\cdots,\eta_{n_{3}})\in(\mathrm{BV}_{0})^{n_{3}}$. We denote $\mathcal{C}^{\mathcal{N}}(\Omega)$ the space of functions admitting continuous such pathwise derivatives of order $\mathcal{N}\subset\mathbb{N}^{3}_{0}$. We can now define PPDEs on $\Omega$ similar to (2.6)
(2.15) $\left\{\hskip-11.38092pt\begin{array}[]{rl}&\mathcal{L}_{t}u(t,x,\gamma)=g(t,x,\gamma),\\ &u(T,x,\gamma)=\phi(T,x,\gamma),\end{array}\right$.
where $g,\phi:\Omega\to\mathbb{R}$ are measurable functions and the linear operator $\mathcal{L}_{t}$ is a combination of derivatives of all arguments, analogue of $\widetilde{\mathcal{L}}_{t}$ (2.7) for functions independent of $\gamma_{[0,t]}$, albeit with a pathwise derivative of Gateaux type instead of Fréchet. More precisely, there exist $\mathcal{N}\subset\mathbb{N}^{3}_{0}$, $a_{n}\in\mathbb{R}$ for all $n\in\mathcal{N}$ and $\bm{\eta}\in(\mathrm{BV}_{0})^{n_{3}}$ such that
@@ 2.5. A kernel method for PPDEs
We saw that the solution to the original PPDE of the type $\widetilde{\mathcal{L}}_{t}u=f$ on $\widetilde{\Omega}$ is a function on continuous paths, whereas signature kernels are defined over paths of bounded variation. For this reason we defined an analogous weaker PPDE $\mathcal{L}_{t}u=f$ on $\Omega\subset\widetilde{\Omega}$ (defined in (2.2),(2.1)) which solution can be suitably represented with signature kernels. Functions admitting Fréchet derivatives also admit Gateaux derivatives and they are equal, therefore solutions to the former PPDE are also solutions to the latter. This will allow us in Theorem 4.5 to bridge the gap between the equation we solve and the equation we target. Now that we have this picture in mind, we can present the gist of our method, inspired by [12] and extended to the path-dependent case.
We consider a kernel $\kappa$, indexed on a compact subset $\mathcal{X}$ of $\Omega$, as product of an RBF kernel and a signature kernel indexed on $\mathbb{R}^{d+1}$ and $\mathrm{BV}_{0}=\mathcal{C}^{1-var}([0,T],\mathbb{R}^{e})$ respectively. We call $\mathcal{H}$ its associated RKHS. We pick $m+n$ collocation points $\omega^{1},...,\omega^{m+n}\in\mathcal{X}$ of the form $\omega^{i}=(t^{i},x^{i},\gamma^{i})$ and order them in such a way that $0\leq t^{i}<T$ for i=1,,mi=1,...,m and $t^{i}=T$ for i=m+1,,m+ni=m+1,...,m+n. The former belong to the interior of the domain while the latter correspond to the boundary (terminal) condition. The main idea consists in approximating the solution of the PPDE (2.15) with a minimiser of the following optimal recovery problem
@@
In other words, we search for the function with minimal RKHS norm that satisfies the PDE constraints at all the collocation points.
This problem formulation raises a number of questions, of both theoretical and practical importance:
@@
    Does the domain of $\mathcal{L}_{t}$ contain $\mathcal{H}$ ?
@@
    Does the PPDE (2.6) have a unique classical solution $u^{\star}$ ?
@@
    Does $u_{m,n}$ converge to $u^{\star}$ as $m,n\to\infty$ ?
After designing our kernel and RKHS in Section 3, we answer positively to all these questions under appropriate conditions. In order to successively execute limit arguments, we exploit that $\mathcal{X}$ is a compact set of $\Omega$, which we define explicitly under the topology induced by $\bm{d}$. The interested reader may find a summary of the answers as follows.
@@
    The optimal recovery problem reduces to a finite-dimensional quadratic optimisation problem with linear constraints. The latter can be solved via gradient descent, aloof from sophisticated learning algorithms. See Section 4.1.
@@
    Theorem 4.5 ensures convergence towards the true solution over a compact subset $\mathcal{X}$ of $\Omega$ under two different sets of assumptions, both requiring that $u^{\star}$ is the unique solution of the PPDE (2.6) and that the collocation points form a countable dense set of $\mathcal{X}$.
PDE Scheme. In particular, after having checked that the constraints of the optimal recovery problem (2.17) make sense, we can solve this optimisation problem explicitly due to a variant of the representer theorem involving higher order derivatives [41], which gives an answer to Question (2) of our list 2.5. Notably, for all $n,m\in\mathbb{N}$ and collocation points $(\omega^{i})_{i=1}^{m}\in\mathcal{X}^{\circ}$, $(\omega^{i})_{i=m+1}^{m+n}\in\partial\mathcal{X}$, there exist weights $\alpha_{1},\cdots,\alpha_{m+n}\in\mathbb{R}$ such that the optimal recovery problem (2.17) is well-posed and has a unique minimiser $u_{m,n}\in\mathcal{H}$ of the form
(2.18) $u_{m,n}=\sum_{i=1}^{m+n}\alpha_{i}\kappa(\omega^{i},\cdot).$
Thus, the coefficients $\alpha_{1},\cdots,\alpha_{m+n}$ can be obtained by simply solving the finite-dimensional quadratic optimisation problem with linear constraints
(2.20) $\widetilde{\mathcal{K}}:=\begin{pmatrix}\mathcal{L}_{t}\kappa(\omega^{1},\omega^{1})&...&\mathcal{L}_{t}\kappa(\omega^{1},\omega^{m+n})\\ \vdots&\ddots&\vdots\\ \mathcal{L}_{t}\kappa(\omega^{m},\omega^{1})&...&\mathcal{L}_{t}\kappa(\omega^{m},\omega^{m+n})\\ \kappa(\omega^{m+1},\omega^{1})&...&\kappa(\omega^{m+1},\omega^{m+n})\\ \vdots&\ddots&\vdots\\ \kappa(\omega^{m+n},\omega^{1})&...&\kappa(\omega^{m+n},\omega^{m+n})\end{pmatrix}\quad\text{and}\quad\bm{b}:=\begin{pmatrix}g(\omega^{1})\\ \vdots\\ g(\omega^{m})\\ \phi(\omega^{m+1})\\ \vdots\\ \phi(\omega^{m+n})\end{pmatrix}$.
Remark 2.2.
For clarity, we are not able at this stage to verify the assumptions of Theorem 4.5 in our examples of interest. We believe however that this paper is the first step towards a more general theory, which shall grow as our understanding of these PPDEs develop. In particular, the assumption that $\mathcal{X}$ is a compact subspace of $\Omega$ (the space of bounded variation paths) could be lifted using at least two strategies we can think of: 1) by carefully rescaling the signature coefficients in a path-by-path exploiting the concept of robust signatures proposed in [14]; or 2) considering a different class of topologies on weighted-spaces as discussed in [20]. Both of these approaches essentially ensure that the range of the signature map remains ”bounded” even when the domain is non-compact. The former does so by changing the feature map, whilst the second changes the underlying path-space topology. At the cost of increased technicality one can also replace the space of bounded variation with a space of arbitrary $p$-variation. Uniqueness of the solution to the PPDE should thus be attainable by classical probabilistic techniques (i.e. Feynman–Kac theorem). For more details about the assumptions we refer to the discussion below Theorem 4.5.
Remark 2.3.
We stress that kernel-based supervised learning approaches like kernel ridge regression require a dataset of input–output pairs, i.e. collocation paths together with corresponding solution values. In our setting such dataset is unavailable: the solution of the PPDE is exactly what we seek to compute. Our solver is unsupervised, in the sense that it enforces the PPDE operator at collocation points as constraint in the kernel-based optimisation in (2.17), without needing prior solution values. Moreover, while conditional expectations can represent PPDE solutions in special cases, for general PPDEs no such probabilistic representation is guaranteed. Our approach therefore applies in broader settings where regression-based methods are not directly applicable.
Remark 2.4.
We note that an alternative approach to our optimal recovery problem would consist in considering instead a “kernel ridge regression”–type formulation, in which one minimises a penalised empirical loss of the PDE residual over the RKHS under a norm constraint or regularisation. A prototypical example is
$\min_{u\in\mathcal{H}}\sum_{i=1}^{m}\big|\mathcal{L}_{t}u(\omega^{i})-g(\omega^{i})\big|^{2}+\sum_{i=m+1}^{m+n}\big|u(\omega^{i})-\phi(\omega^{i})\big|^{2}+\lambda\|u\|_{\mathcal{H}}^{2},$
for some regularisation parameter $\lambda>0$, or more generally a formulation based on an $L^{2}$–loss of the PDE residual over a measure on the domain. Results of this type can be found, for instance, in [57, Theorem 2]. Such kernel ridge regression (KRR) discretisation based on the same kernel ansatz
$u(\cdot)\;=\;\sum_{j=1}^{m+n}\alpha_{j}\,\kappa(\cdot,\omega^{j})$
leads to the unconstrained penalised least-squares problem
$\min_{\bm{\alpha}\in\mathbb{R}^{m+n}}\;\big\|\widetilde{\mathcal{K}}\,\bm{\alpha}-\bm{b}\big\|_{2}^{2}\;+\;\lambda\,\bm{\alpha}^{\top}\mathcal{K}\bm{\alpha},\qquad\lambda>0,$
where the same matrices $\mathcal{K}$, $\widetilde{\mathcal{K}}$ and right-hand side $\bm{b}$ appear. The first term measures the squared $\ell^{2}$-residual of the PDE and boundary/terminal constraints at the collocation points, while the second term penalises the RKHS norm of $u$ via
$\|u\|_{\mathcal{H}}^{2}=\bm{\alpha}^{\top}\mathcal{K}\bm{\alpha}.$
The associated first-order optimality conditions yield the normal equations
$\big(\widetilde{\mathcal{K}}^{\top}\widetilde{\mathcal{K}}+\lambda\mathcal{K}\big)\bm{\alpha}\;=\;\widetilde{\mathcal{K}}^{\top}\bm{b}.$
Thus, both discretisations are built from the same kernel expansion and the same matrices $(\mathcal{K},\widetilde{\mathcal{K}},\bm{b})$. The only structural difference is that the optimal recovery formulation imposes
$\widetilde{\mathcal{K}}\,\bm{\alpha}=\bm{b}$
as hard constraints and minimises the RKHS norm subject to these constraints, whereas kernel ridge regression replaces these constraints by a quadratic penalty and enforces them only in a least-squares sense. In particular, for any fixed $\lambda>0$ one generally has
$\widetilde{\mathcal{K}}\,\bm{\alpha}\neq\bm{b},$
so that the PDE and boundary/terminal conditions are not satisfied exactly at the collocation points, with the trade-off between fidelity to the constraints and RKHS norm controlled by the regularisation parameter $\lambda$. Formally, the optimal recovery problem may be viewed as the zero-regularisation or infinite-penalty limit of kernel ridge regression: as $\lambda\downarrow 0$, any accumulation point of minimisers of the penalised problem that satisfies $\widetilde{\mathcal{K}}\,\bm{\alpha}=\bm{b}$ coincides with the minimum-norm solution of the constrained problem. Conversely, for fixed $\lambda>0$, the kernel ridge solution does not in general satisfy the PDE exactly at the collocation points.
@@
Finally, let us summarise the several levels of approximation. We numerically solve the optimisation problem (2.17), which is a discretisation of the PPDE (2.15) over $\Omega$. Under the right set of assumptions, the sequence of solutions to the optimisation problem converges the solution to the PPDE (2.6) over $\widetilde{\Omega}$. Finally, the latter is an approximation of the PPDE with a singular direction $\bm{\eta}$. While our setup is designed to match PPDEs of second generation (arising from Volterra processes), it could be adapted to the first generation of Dupire (corresponding to path-dependent payoffs) at the cost of redeveloping the theory with the corresponding spaces, notions of derivatives and topologies.